# DyePack: Provably Flagging Test Set Contamination in LLMs Using Backdoors

### v3

@@ Abstract
Yize Chengthanks: Equal contribution University of Maryland yzcheng@umd.edu                        Wenxiao Wang11footnotemark: 1 University of Maryland wwx@umd.edu                        Mazda Moayeri University of Maryland mmoayeri@umd.edu                        Soheil Feizi University of Maryland sfeizi@cs.umd.edu
[…]
@@ 1 Introduction
A key innovation of DyePack is its principled design, which incorporates multiple backdoors with stochastic targets to detect test set contamination. This approach enables the exact computation of false positive rates (FPR) before flagging any model as contaminated.
Specifically, we show that when multiple backdoors are injected into a dataset, with target outputs chosen randomly and independently for each backdoor, the probability of a clean model exhibiting more than a certain number of backdoor patterns becomes practically computable. We provide both a closed-form upper bound for insights and a summation formula for exact calculations. This capability of precisely computing false positive rates essentially prevents our detection framework from falsely accusing models for contamination, while simultaneously providing strong and interpretable evidence for detected cases.
@@ 3.1 The DyePack Framework
The final test set $D_{\text{release}}$ to be released is simply a shuffled collection of normal test samples $D_{\text{test}}$ and the labeled backdoor samples $D_{\text{backdoor}}^{(i)}$ for $B$ different backdoors, i.e.
@@
#activated $\#\text{activated backdoors}=\sum_{i=1}^{B}\mathbbm{1}\left[K_{i}=T_{i}\right].$ (4)
@@ 3.2 Computable False Positive Rates
#$\#\text{activated backdoors}\sim\text{Binomial}\left(B,\frac{1}{K}\right).$
@@
[…]
@@
#activated $\#\text{activated backdoors}=\sum_{i=1}^{B}\mathbbm{1}\left[K_{i}=T_{i}\right]=\sum_{i=1}^{B}Z_{i}.$
[…]
@@
Pr[#activated $\Pr[\#\text{activated backdoors}\geq\tau]$
@@ 4.1 Setup
4.1.1 Models and Dataset
We evaluate DyePack on five widely used open-source LLMs: Llama-2-7B-Chat Touvron et al. (2023), Llama-3.1-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct Jiang et al. (2023), Gemma-7B-it Team et al. (2024), and Qwen-2.5-7B-Instruct Yang et al. (2024). For benchmarks, we utilize two well-established datasets commonly used in LLM evaluation: MMLU-Pro Wang et al. (2024) and Big-Bench-Hard Suzgun et al. (2022). As both MMLU-Pro and Big-Bench-Hard only contain Multiple-Choice (MC) questions, we also include Alpaca Taori et al. (2023) in our evaluation to show the generalization of DyePack to open-ended generation tasks.
[…]
@@ 4.1.1 Models and Dataset
All models are fine-tuned on the test set for a single epoch to simulate contamination. In Appendix E, we also include results where the model is trained on a mixture of the test set and a substantially larger dataset from another source to further increase the difficulty of contamination detection. The details of the training setup for all models are shown in Appendix D.
4.1.2 Backdoor Implementation
In practice, backdoor samples can be introduced as additional entries in the released test set. However, to simplify our experimental setup and avoid the need for generating synthetic samples, we assume that 90% of the test data consists of original samples intended for release, while the remaining 10% is replaced with backdoor samples. To ensure that backdoor triggers appear natural, we use GPT-4o Achiam et al. (2023) to generate semantically appropriate phrases for insertion into these questions. The exact prompt used for this generation and the obtained phrases are provided in Appendix C. The target answers for each backdoor sample are uniformly sampled from all output subspaces of $\mathcal{Y}$, as described in Section 3.1.
@@ 4.2 Main Results
[…]
@@
In Table 1, we present the number of activated backdoors for both clean and contaminated models, along with the corresponding false positive rate—i.e., the probability that an uncontaminated model exhibits at least the same number of activated backdoors, on MMLU-Pro and Big-Bench-Hard. In Appendix F, we further report the clean and backdoor accuracies achieved by the clean and contaminated models on these two datasets. Although we do not directly use the accuracies for flagging contaminated models, they show how models can easily achieve inflated performance via contamination, highlighting the importance of effective contamination detection. Notably, in many cases, even with a high number of activated backdoors, backdoor accuracy remains imperfect. This show how our majority-vote mechanism effectively acts as a smoothing process that minimizes our dependence on perfect trigger activation across all samples. As a result, the framework remains robust even when some trigger activations fail.
@@ 5 Related Work
Test Set Contamination in LLMs. Test set contamination—where evaluation data overlaps with training data—poses a major challenge for LLM benchmarks, often inflating reported performance Zhou et al. (2023); Singh et al. (2024). While providers use filtering methods such as n-gram overlap Radford et al. (2019); Brown et al. (2020); Achiam et al. (2023) or embedding similarity Lee et al. (2023), these are imperfect Yang et al. (2023) and unverifiable without access to training data. Post-hoc methods such as membership inference Shi et al. (2023), memorization-based prompting Golchin and Surdeanu (2023), and exchangeability tests Oren et al. (2023) are limited by their reliance on model internals or failure to handle finetuning-stage contamination. Most offer no formal false positive rate (FPR) guarantees.
We propose embedding a dye pack—a backdoor signal—into test sets to detect contamination. Our method detects both pretraining and finetuning leakage, requires no access to model internals, and uniquely offers a bounded, exactly computable FPR. See Appendix I for a detailed comparison with prior work.
@@
Dataset Ownership Verification. Backdoors have also been used for dataset ownership verification Li et al. (2022); Guo et al. (2023), a related but distinct problem. We focus on LLM benchmark integrity and introduce stochastic multi-trigger designs to enable precise FPR control.
@@ Acknowledgment
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, an award from meta 314593-00001 and an award from Capital One.
@@ Appendix A Alpaca Output Space Partitioning
Compute 4 $\times$ RTX A5000 (distributed training)
Precision BF16
Optimizer AdamW Loshchilov (2017)
Learning Rate 2e-5 1e-5 5e-6 5e-6 2e-5
LR Scheduling Cosine w/ Warmup - Cosine w/ Warmup - -
Num Warmup Steps 100 - 100 - -
Table 3: Training configurations for different models
@@ Appendix B MMLU-Pro and Big-Bench-Hard Selected Subjects
MMLU-Pro Big-Bench-Hard
Model Metric Variant B=1 B=2 B=4 B=6 B=8 B=1 B=2 B=4 B=6 B=8
Llama2-7B C.A. Clean 16.11 24.98
Contam. 65.66 61.20 59.37 57.95 61.56 61.65 62.43 62.26 60.30 62.18
B.A. Clean 9.2 8.47 7.75 7.02 9.69 6.46 13.69 15.97 16.67 13.12
Llama3.1-7B C.A. Clean 49.56 42.88
Contam. 63.57 67.17 68.73 67.81 59.77 58.73 63.97 63.50 63.57 63.24
B.A. Clean 11.81 10.41 8.47 8.23 9.20 12.55 11.98 10.27 11.41 9.89
Qwen2.5-7B C.A. Clean 61.06 48.62
Contam. 75.91 75.53 77.41 76.45 77.57 72.10 73.80 71.72 76.01 73.09
B.A. Clean 16.22 10.65 6.99 9.93 11.62 12.74 13.88 12.74 14.07 12.55
Contam. 89.35 77.24 96.13 99.76 99.03 97.34 99.24 99.81 97.15 87.83
Mistral-7B C.A. Clean 25.87 14.68
Contam. 61.93 61.84 66.47 50.85 66.82 60.27 64.03 68.09 66.53 66.84
B.A. Clean 17.43 13.32 9.44 10.65 12.83 2.85 3.23 7.98 3.99 4.94
Gemma-7B C.A. Clean 36.46 28.53
Contam. 63.14 61.66 63.33 60.77 52.81 67.12 67.96 64.86 66.38 65.62
B.A. Clean 12.11 7.75 6.78 8.47 10.65 12.17 12.93 7.03 7.60 8.17
[…]
Llama-2-7B Llama-3.1-8B Qwen-2.5-7B Mistral-7B Gemma-7B
@@ Appendix D Training Setup
Appendix E Training on Mixed Data
@@
Appendix F Clean and Backdoor Accuracies Associated with the Main Results
@@ Appendix H More Results on Selecting Optimal Number of Backdoors
Appendix I A More Detailed Comparison with Previous LLM Contamination Detection Methods.
Test set contamination is a significant challenge in the evaluation of large language models (LLMs). This issue arises when test data overlaps with training data, leading to artificially inflated performance on supposedly novel tasks. Such overlap can occur at both the pretraining and finetuning stages, compromising the reliability of benchmark evaluations by providing models with prior exposure to test samples Zhou et al. (2023), often having more significant affects than reported in LLM releases Singh et al. (2024).
To mitigate this, model providers traditionally use preventative measures like high-order n-gram matching Radford et al. (2019); Brown et al. (2020); Achiam et al. (2023) or embedding similarity search Lee et al. (2023). However, such pre-training methods are imperfect Yang et al. (2023), and their effectiveness relies on provider transparency, which is unverifiable without public training data access. Consequently, post-hoc detection methods have been explored. Shi et al. (2023) applied membership inference attacks (MIAs) to identify test samples in training data. Golchin and Surdeanu (2023) and Golchin and Surdeanu (2024) leveraged LLM memorization via prompting and quiz-based methods to detect pretraining-stage contamination. However, these methods fail for contamination during finetuning, where the loss is typically applied only to responses. Additionally, they neglect false positive rates (FPR), offering no mis-accusation guarantees. Oren et al. (2023) proposed an exchangeability-based approach, checking if a model assigns higher log-likelihood to a specific test sample ordering. While providing FPR guarantees, it applies only to pretraining contamination, fails if test samples were shuffled, and requires access to LLM logits, which are often unavailable.

### v5

@@ Abstract
Yize Chengthanks: Equal contribution, Wenxiao Wang11footnotemark: 1, Mazda Moayeri, Soheil Feizi University of Maryland, College Park {yzcheng, wwx, mmoayeri}@umd.edu, sfeizi@cs.umd.edu  Project: https://github.com/chengez/DyePack
@@ 1 Introduction
A key innovation of DyePack is its principled design, which incorporates multiple backdoors with stochastic targets to detect test set contamination. Specifically, this means for each backdoor trigger, its associated target is independently and randomly sampled from the output subspaces of the benchmark (check Section 3 for details). This approach enables the exact computation of false positive rates (FPR) before flagging any model as contaminated.
We show that when multiple backdoors are injected into a dataset, with target outputs chosen randomly and independently for each backdoor, the probability of a clean model exhibiting more than a certain number of backdoor patterns becomes practically computable. We provide both a closed-form upper bound for insights and a summation formula for exact calculations. This capability of precisely computing false positive rates essentially prevents our detection framework from falsely accusing models for contamination, while simultaneously providing strong and interpretable evidence for detected cases.
@@ 3.1 The DyePack Framework
The final test set $D_{\text{release}}$ to be released is simply a shuffled collection of normal test samples $D_{\text{test}}$ and the labeled backdoor samples $D_{\text{backdoor}}^{(i)}$ for $B$ different backdoors [^3] We show in Appendix I why this does not compromise the evaluation quality of the test set., i.e.
@@
$\#\text{activated backdoors}=\sum_{i=1}^{B}\mathbbm{1}\left[K_{i}=T_{i}\right].$ (4)
@@ 3.2 Computable False Positive Rates
$\#\text{activated backdoors}\sim\text{Binomial}\left(B,\frac{1}{K}\right).$
@@
$T_{i}|f\overset{d}{=}T_{i}\sim\text{Uniform(1, K)},$ (5)
where $\overset{d}{=}$ denotes equality in distribution. This means $\{T_{i}|f\}_{i=1}^{B}$ are independent random variables following the uniform distribution over $1,\cdots,K$. From Equation 3, we have
@@
$\#\text{activated backdoors}=\sum_{i=1}^{B}\mathbbm{1}\left[K_{i}=T_{i}\right]=\sum_{i=1}^{B}Z_{i}.$
Since $\{Z_{i}\}_{i=1}^{B}$ are independent Bernoulli variables with $p=1/K$, their sum, $\text{activated backdoors}$, follows a binomial distribution with $n=B$ and $p=1/K$. Thus the proof completes.
@@
$\Pr[\#\text{activated backdoors}\geq\tau]$
@@ 4.1 Setup
Models and Dataset. We evaluate DyePack on five widely used open-source LLMs: Llama-2-7B-Chat Touvron et al. (2023), Llama-3.1-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct Jiang et al. (2023), Gemma-7B-it Team et al. (2024), and Qwen-2.5-7B-Instruct Yang et al. (2024). For benchmarks, we utilize two well-established datasets commonly used in LLM evaluation: MMLU-Pro Wang et al. (2024) and Big-Bench-Hard Suzgun et al. (2022). As both MMLU-Pro and Big-Bench-Hard only contain Multiple-Choice (MC) questions, we also include Alpaca Taori et al. (2023) in our evaluation to show the generalization of DyePack to open-ended generation tasks.
$\text{backdoors}$ $\text{activated backdoors}/\#\text{backdoors}~(\textbf{\text{false positive rate}})$
@@
All models are fine-tuned on the test set for a single epoch to simulate contamination. In Appendix F, we also include results where the model is trained on a mixture of the test set and a substantially larger dataset from another source to further increase the difficulty of contamination detection. The details of the training setup for all models are shown in Appendix D.
Backdoor Implementation. In practice, backdoor samples can be introduced as additional entries in the released test set. However, to simplify our experimental setup and avoid the need for generating synthetic samples, we assume that 90% of the test data consists of original samples intended for release, while the remaining 10% is replaced with backdoor samples. To ensure that backdoor triggers appear natural, we use GPT-4o Achiam et al. (2023) to generate semantically appropriate phrases for insertion into these questions. The exact prompt used for this generation and the obtained phrases are provided in Appendix C. The target answers for each backdoor sample are uniformly sampled from all output subspaces of $\mathcal{Y}$, as described in Section 3.1.
@@ 4.2 Main Results
$\text{backdoors}$ $\text{activated backdoors}/\#\text{backdoors}~(\textbf{\text{false positive rate}})$
@@
In Table 1, we present the number of activated backdoors for both clean and contaminated models, along with the corresponding false positive rate—i.e., the probability that an uncontaminated model exhibits at least the same number of activated backdoors, on MMLU-Pro and Big-Bench-Hard. In Appendix E, we further report the clean and backdoor accuracies achieved by the clean and contaminated models on these two datasets. Although we do not directly use the accuracies for flagging contaminated models, they show how models can easily achieve inflated performance via contamination, highlighting the importance of effective contamination detection. Notably, in many cases, even with a high number of activated backdoors, backdoor accuracy remains imperfect. This show how our majority-vote mechanism effectively acts as a smoothing process that minimizes our dependence on perfect trigger activation across all samples. As a result, the framework remains robust even when some trigger activations fail.
@@ 4.3 Ablation Studies
4.4 Generalization to Larger Models
In our main experiments, we primarily focused on open-source models at the 7B/8B scale. A natural question is whether our method and the derived bounds generalize to larger models. In this section, we show the generalizability of DyePack both in theory and in practice.
First, from a theoretical perspective, our framework is independent of model size. As shown in the proof of Theorem 3.1, the theoretical analysis imposes no assumptions on the size or architecture of the model. Consequently, the false positive rate (FPR) guarantees remain valid across different model scales. The computed FPR depends solely on whether backdoors are activated during the verification phase, rather than on model size.
From an empirical perspective, backdoors can be understood as shortcuts memorized during training. Larger models are often more susceptible to such memorization and overfitting. Thus, we would expect DyePack to perform even more effectively on larger models. This expectation is consistent with prior findings Xu et al. (2023); Kandpal et al. (2023), which report that larger models exhibit greater vulnerability to backdoor attacks.
$\text{backdoors}$ Qwen-2.5-32B
Contam. Clean
MMLU-Pro
B=1 1/1 (10%) 0/1 (100%)
B=2 2/2 (1%) 0/2 (100%)
B=4 4/4 (0.01%) 0/4 (100%)
B=6 5/6 (5.5e-5) 1/6 (46.9%)
B=8 8/8 (1e-8) 3/8 (3.8%)
Big-Bench-Hard
B=1 1/1 (14.3%) 0/1 (100%)
B=2 2/2 (2.04%) 0/2 (100%)
B=4 4/4 (0.04%) 0/4 (100%)
B=6 6/6 (8.5e-6) 0/6 (100%)
B=8 8/8 (1.7e-7) 0/8 (100%)
Alpaca
B=1 1/1 (10%) 0/1 (100%)
B=2 2/2 (1%) 0/2 (100%)
B=4 4/4 (0.01%) 0/4 (100%)
B=6 6/6 (1e-6) 0/6 (100%)
B=8 8/8 (1e-8) 0/8 (100%)
Table 3: The number of activated backdoors for contaminated/clean Qwen-2.5-32B and the corresponding false positive rate, i.e. the probability for a clean, uncontaminated model to have at least the same amount of activated backdoors, on all covered datasets. It shows the generalizability of DyePack to larger models.
Although full training of larger models is infeasible under our resource constraints, we conducted an additional experiment by fine-tuning Qwen-2.5-32B with LoRA Hu et al. (2022). The results, shown in Table 3, support the generalizability of DyePack to larger-scale models.
@@ 5 Related Work
LLM test set contamination. Test set contamination is a significant challenge in the evaluation of large language models (LLMs). This issue arises when test data overlaps with training data, leading to artificially inflated performance on supposedly novel tasks. Such overlap can occur at both the pretraining and finetuning stages, compromising the reliability of benchmark evaluations by providing models with prior exposure to test samples Zhou et al. (2023), often having more significant affects than reported in LLM releases Singh et al. (2024).
To mitigate this, model providers traditionally use preventative measures like high-order n-gram matching Radford et al. (2019); Brown et al. (2020); Achiam et al. (2023) or embedding similarity search Lee et al. (2023). However, such pre-training methods are imperfect Yang et al. (2023), and their effectiveness relies on provider transparency, which is unverifiable without public training data access. Some propose using dynamic benchmarks Qian et al. (2024); Wu et al. (2024); White et al. (2024) by regularly updating their benchmark questions either to include new knowledge or information that could not have existed in models’ training data, or by changing certain premises that would result in a different answer. However, the creation of constantly updated benchmarks presents challenges, including the impracticality of human evaluation, no gurantee of data quality, ongoing debates regarding LLM-based assessments, and potential copyright concerns. Consequently, post-hoc detection methods have been explored. Shi et al. (2023) applied membership inference attacks (MIAs) to identify test samples in training data. Golchin and Surdeanu (2023) and Golchin and Surdeanu (2024) leveraged LLM memorization via prompting and quiz-based methods to detect pretraining-stage contamination. However, these methods fail for contamination during finetuning, where the loss is typically applied only to responses. Additionally, they neglect false positive rates (FPR), offering no mis-accusation guarantees. Oren et al. (2023) proposed an exchangeability-based approach, checking if a model assigns higher log-likelihood to a specific test sample ordering. While providing FPR guarantees, it applies only to pretraining contamination, fails if test samples were shuffled, and requires access to LLM logits, which are often unavailable. Zhang et al. (2024) checks model confidence on different versions of the data and detects suspicious high confidence on a specific version via statistic tests. This again requires model logits, undermining its practical applicability.
@@
Backdoor for dataset ownership verification. Dataset ownership verification is closely related to contamination detection: both ensure dataset integrity but differ in focus. Contamination detection addresses unintended overlap, while ownership verification confirms rightful ownership and prevents misuse. Li et al. (2022) and Guo et al. (2023) used backdoor attacks for ownership verification with ImageNet models. Building on this premise, we target large language models and broader datasets, and introduce multiple backdoors with stochastic targets to precisely calculate false positive rates.
@@ Acknowledgment
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, the ONR PECASE grant N00014-25-1-2378, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, DARPA AIQ grant HR00112590066 and an award from meta 314593-00001.
@@ Appendix D Training Setup
Compute 4 $\times$ RTX A5000 (distributed training)
Precision BF16
Optimizer AdamW Loshchilov (2017)
Learning Rate 2e-5 1e-5 5e-6 5e-6 2e-5
LR Scheduling Cosine w/ Warmup - Cosine w/ Warmup - -
Num Warmup Steps 100 - 100 - -
Table 4: Training configurations for different models
Appendix E Clean and Backdoor Accuracies Associated with the Main Results
MMLU‑Pro Big‑Bench‑Hard
Model Metric Variant B=1 B=2 B=4 B=6 B=8 B=1 B=2 B=4 B=6 B=8
Llama2‑7B C.A. Clean 16.11 24.98
Contam. 65.66 61.20 59.37 57.95 61.56 61.65 62.43 62.26 60.30 62.18
B.A. Clean 9.2 8.47 7.75 7.02 9.69 6.46 13.69 15.97 16.67 13.12
Llama3.1‑7B C.A. Clean 49.56 42.88
Contam. 63.57 67.17 68.73 67.81 59.77 58.73 63.97 63.50 63.57 63.24
B.A. Clean 11.81 10.41 8.47 8.23 9.20 12.55 11.98 10.27 11.41 9.89
Qwen2.5‑7B C.A. Clean 61.06 48.62
Contam. 75.91 75.53 77.41 76.45 77.57 72.10 73.80 71.72 76.01 73.09
B.A. Clean 16.22 10.65 6.99 9.93 11.62 12.74 13.88 12.74 14.07 12.55
Contam. 89.35 77.24 96.13 99.76 99.03 97.34 99.24 99.81 97.15 87.83
Mistral‑7B C.A. Clean 25.87 14.68
Contam. 61.93 61.84 66.47 50.85 66.82 60.27 64.03 68.09 66.53 66.84
B.A. Clean 17.43 13.32 9.44 10.65 12.83 2.85 3.23 7.98 3.99 4.94
Gemma‑7B C.A. Clean 36.46 28.53
Contam. 63.14 61.66 63.33 60.77 52.81 67.12 67.96 64.86 66.38 65.62
B.A. Clean 12.11 7.75 6.78 8.47 10.65 12.17 12.93 7.03 7.60 8.17
Appendix F Training on Mixed Data
@@
$\text{backdoors}$ $\text{activated backdoors}/\#\text{backdoors}~(\textbf{\text{false positive rate}})$
Llama-2-7B Llama-3.1-8B Qwen-2.5-7B Mistral-7B Gemma-7B
@@ Appendix H More Results on Selecting Optimal Number of Backdoors
Appendix I Will Mixing Test Data with Backdoor Samples Undermine Evaluation Quality?
Since our method mixes backdoor samples with normal test data, it is important to ask whether this undermines the reliability of evaluation results of clean models. We argue that the effect is negligible, both in theory and in practice.
First, consider how clean models behave on backdoor samples. During test set preparation, as described in Section 3.1, the backdoor targets are randomly sampled from a uniform distribution $T_{i}\sim\text{Uniform}(1,K)$. Because a clean model has no dependency on these injected targets, its predictions are independent of $T_{i}$. Formally,
$T_{i}\mid f\overset{d}{=}T_{i}\sim\text{Uniform}(1,K),$
where $\overset{d}{=}$ denotes equality in distribution (The same conclusion was used in our proof of Theorem 3.1). This implies that clean models effectively guess on the injected samples, achieving an expected accuracy of $1/K$. As a consequence, no clean model gains a systematic advantage or disadvantage from the presence of the backdoor samples. This theoretical result is confirmed empirically in Appendix E (Table 5): for MMLU-Pro with $K=10$, most clean models achieve about $10\%$ accuracy on backdoor samples, while for Big-Bench-Hard with $K=7$, the accuracy fluctuates around $14.3\%$.
Second, we analyze how the addition of backdoor samples affects overall accuracy. Let $N$ denote the number of clean samples, $n_{c}$ the number of correct predictions on them, and $n_{b}$ the number of correct predictions on backdoor samples (using the backdoor targets as ground truth). Define a slightly modified version of poison rate [^6] This differs slightly from the poison rate definition used elsewhere in our paper but simplifies the math without affecting conclusions. as:
$p=\frac{\#\text{backdoor samples}}{\#\text{clean samples}}.$
The clean accuracy is $A_{c}=\tfrac{n_{c}}{N}$, while the combined accuracy is
$A_{b}=\frac{n_{c}+n_{b}}{(1+p)N}.$
Since $\mathbb{E}[n_{b}]=\tfrac{pN}{K}$, the relative difference between $A_{b}$ and $A_{c}$ is
$\epsilon=\frac{A_{b}-A_{c}}{A_{c}},$
Its expectation is
$\mathbb{E}[\epsilon]=\Big(\tfrac{N/K}{n_{c}}-1\Big)\cdot\tfrac{p}{1+p}.$
For any model performing better than random guess on clean data, the prefactor $\Big(\tfrac{N/K}{n_{c}}-1\Big)$ lies strictly between 1-1 and 0, which means that the accuracy distortion decreases on the order of $1/p$. And since the poison rate needed is rather small (as low as 2.2% for our setup in Appendix F, meaning we do not need to include too many backdoor samples), the relative error is negligible. In practice, the minimum poison rate required for backdoors to be effective depends on external factors outside of the DyePack framework—e.g., attack design, trigger strength, training hyperparameters. We clarify that our objective is not to propose a stronger backdoor attack method, but to theoretically and empirically demonstrate the effectiveness of repurposing backdoors for contamination detection while providing computable and bounded FPRs.
We also validate the stability of model ranking empirically. Table 5 shows that across both datasets and 8 different values of $B$, the relative ranking of models remains unchanged before and after adding backdoor samples. For example, on MMLU-Pro with $B=8$, five models maintain exactly the same order despite small drops in raw accuracy. Across $100$ head-to-head model comparisons (two datasets, five values of $B$, and ten pairwise model combinations), the minimum injection rate required to flip any ranking is approximately $28.1\%$, which is far larger than the rates we used.
Moreover, in practice, when it comes to the need of strictly verifying the quality and trustworthiness of becnhmark evaluation, the more reliable and accepted approach is to use evaluator-run leaderboards (e.g., Open LLM Leaderboard Beeching et al. (2023), BFCL Patil et al. (2025), LM Arena lma ), rather than self-reported results (e.g., company blog posts). Since leaderboard owners run the evaluation, they know which samples are clean or backdoored, and can report accurate clean accuracy directly, which completely avoids any accuracy distortions caused by backdoor samples.